OpenStudy (anonymous):

Let f (x) = x2 + 18x + 85. (a) Find the vertex. Answer: ________ (b) State the range of the function. Answer: ________ (c) On what interval is the function decreasing? Answer: ________

OpenStudy (anonymous):

f(x)=\[x ^{2}+18x+85\] f(x)=\[(x+9)^2-81+85\] \[f(x)=(x+9)^2+4\] Vertex = (-9 , 4) \[range = [4.\infty[\] It decreases \[x \in ]-\infty,-9]\]

OpenStudy (anonymous):

Very good!